OpenStudy (anonymous):

Estimate 3 and two-sevenths times 4 and three-fourths by rounding each factor to the nearest whole number. Explain your choices for rounding

OpenStudy (anonymous):

Is (2/7) enough to turn the 3 into a 4? Is (3/4) enough to turn the 4 into a 5?

OpenStudy (unklerhaukus):

\[(3+2/7) \times ( 4+3/4)\]\[\approx(3+0)\times(4+1)\]\[=3\times5\]